# Less is More: Compact Clue Selection for Efficient Retrieval-Augmented Generation Reasoning

### v6

@@ Appendix C More Experimental Results on LLaMA3-70B
Table 12. Experimental results on three benchmark datasets using LLaMA3.3-70B-Instruct as the generator. The retriever is DPR. The experimental setup follows Section 5.1.4. All baselines and our method are conducted using the same test sets and retrieval corpus. SubEM = substring Exact Match, F1 = F1 score, CR = $\frac{\#\text{tokens in retrieved documents}}{\#\text{tokens in compressed text}}$. The best results are in bold and the second are underlined.
@@ C.2. Comparison with Baselines
Following the same experimental setting described in Section 5.1.4, we conducted experiments on NQ, TriviaQA, and HotpotQA. We use the DPR retriever to retrieve the Top-5 documents and adopt LLaMA3.3-70B-Instruct as the generator. All baselines and our method are conducted using the same test sets and retrieval corpus. The results are summarized in Table C. Across all three datasets, our method consistently outperforms various baseline approaches. This demonstrates that our approach effectively selects the most informative context from retrieved documents, enhancing performance on downstream generation tasks. Moreover, the consistent performance gains highlight the stability and strong generalization ability of our method across different task distributions, making it applicable in a wide range of retrieval and generation scenarios.
@@ C.3. Comparison with Full Content Strategy
This section evaluates the generalization ability of our method across different retrievers with all experiments conducted on the NQ dataset. Specifically, we train the model using documents retrieved by the DPR retriever and perform inference on documents retrieved by the Contriever-MS MARCO retriever, simulating a cross-retrieval distribution shift. Two extractors trained with LLaMA3.2-3B-Instruct (described as 3B Extractor) and LLaMA3-8B-Instruct (described as 8B Extractor) are employed, respectively, and compared against the Full Content baseline. To comprehensively assess performance under varying context sizes, we select representative settings from Top-1 to Top-60 retrieved documents for comparison, allowing us to understand how context scale impacts model performance and the benefits of selective extraction.
@@
Additionally, increasing the size of the extractor from 3B to 8B generally improves downstream task performance, although it typically results in higher inference latency. This implies a trade-off between performance and efficiency in practical applications. Larger extractors are able to more accurately select key information, enhancing generation quality under large-scale contexts, but at the cost of increased computational resources and response time. In scenarios with limited resources or stringent real-time requirements, choosing an appropriately sized extractor is crucial, whereas in research settings or batch processing tasks aiming for the highest generation performance, employing a larger extractor can offer clear advantages.

### v7

@@ Appendix C More Experimental Results on LLaMA3-70B
Table 12. Experimental results (%) on three benchmark datasets using LLaMA3.3-70B-Instruct as the generator. The retriever is DPR. The experimental setup follows Section 5.1.4. All baselines and our method are conducted using the same test sets and retrieval corpus.
@@ C.2. Comparison with Baselines
Following the same experimental setting described in Section 5.1.4, we conduct experiments on NQ, TriviaQA, and HotpotQA. We use the DPR retriever to retrieve the Top-5 documents and adopt LLaMA3.3-70B-Instruct as the generator. All baselines and our method are conducted using the same test sets and retrieval corpus. The results are summarized in Table C. Across all three datasets, our method consistently outperforms various baseline approaches. This demonstrates that our approach effectively selects the most informative context from retrieved documents, enhancing performance on downstream generation tasks. Moreover, the consistent performance gains highlight the stability and strong generalization ability of our method across different task distributions, making it applicable in a wide range of retrieval and generation scenarios.
@@ C.3. Comparison with Full Content Strategy
This section evaluates the generalization of our method across retrievers on the NQ dataset. We train with documents from the DPR retriever and test on the Contriever-MS MARCO retriever, simulating cross-retrieval shifts. Two extractors, LLaMA3.2-3B-Instruct (3B Extractor) and LLaMA3-8B-Instruct (8B Extractor), are compared against the Full Content baseline across Top-1 to Top-60 retrieved documents to analyze context-scale effects.
@@
Increasing the extractor size from 3B to 8B improves performance but increases latency, highlighting a trade-off between efficiency and quality. Larger extractors better select key information for large-scale contexts, whereas smaller extractors may be preferable under limited resources or real-time constraints.
@@
Appendix D Case Study
As shown in Figure 7, our method highlights potential clues (red background), reranks them, and surfaces the correct answer clue at the top, while filtering out redundant ones to improve the information density of reasoning clues for RAG.